OpenStudy (anonymous):

Solve the problem, If f(x)= x-3A/3x+5 and f(3)=6. what is the value of A?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

this is f(x), right? \[\Large f(x) = \frac{x-3A}{3x+5}\]

OpenStudy (anonymous):

Correct

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{x-3A}{3x+5}\] \[\Large f(3) = \frac{3-3A}{3(3)+5}\] \[\Large 6 = \frac{3-3A}{3(3)+5}\] \[\Large 6 = \frac{3-3A}{14}\] \[\Large 14*6 = 14*\frac{3-3A}{14}\] \[\Large 84 = 3-3A\] \[\Large A = ??\]

jimthompson5910 (jim_thompson5910):

In step 2, I replaced every x with 3 In step 3, I replaced f(3) with 6 (since f(3) = 6)

OpenStudy (anonymous):

Okay so I got A= -27

OpenStudy (anonymous):